OpenStudy (anonymous):

Show that the equation 2tan^2 theta cos theta= 3 can be written in the form 2cos^2 theta + 3 cos theta=0

OpenStudy (stacey):

\[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\]Use substitution I think you have an error in your problem. I end up with \[2\cos ^{2}\theta+3\cos \theta = 2\]